# On-Policy Self-Alignment with Fine-grained Knowledge Feedback for Hallucination Mitigation

### v5

@@ Abstract
Xueru Wen1,2, Jie Lou3, Xinyu Lu1,2, Ji Yuqiu3, Xinyan Guan1,2, Yaojie Lu1, Hongyu Lin1, Ben He1,2, Xianpei Han1,2, Debing Zhang3, Le Sun1,2 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences, Beijing, China 2University of Chinese Academy of Sciences, Beijing, China 3Xiaohongshu Inc {wenxueru2022,luxinyu2021,guanxinyan2022}@iscas.ac.cn {luyaojie,hongyu,sunle,xianpei}@iscas.ac.cn benhe@ucas.edu.cn loujie0822@gmail.com dengyang@xiaohongshu.com
@@ 1 Introduction
    Comprehensive evaluations on HotpotQA, SQuADv2, and Biography present significant improvements of RLFH over both base models and existing hallucination mitigation approaches, showing the method’s effectiveness.
@@ 3.2.2 Online Reinforcement Learning
Specifically, we first sample the prompt $x$ and corresponding response $y$. Subsequently, the policy model $\pi$ itself serves as the judge to provide fine-grained feedback through the assessment framework. This feedback is then converted into token-level dense reward $r=\left[r_{1},r_{2},...,r_{T}\right]$, where $T$ denotes the total length of the response $y$. Finally, we employ this reward $r$ to update the policy model $\pi$ by the Proximal Policy Optimization (Schulman et al., 2017a) algorithm.
@@ 4.2 Main Results
Table 2: Results of baseline and RLHF-trained models using different granularity reward signals.
@@ Appendix D Factuality Metrics Validation
Appendix E KL Estimation
In this work, we use Proximal Policy Optimization (PPO) Schulman et al. (2017b) to learn policies. PPO introduces a trust region mechanism that restricts policy updates to prevent performance collapse during training. This trust region is implemented by constraining the Kullback-Leibler (KL) divergence between the new policy and the old policy, which requires efficient KL estimation during the training process.
There are three typical approaches to estimate KL divergence (Schulman, 2020). Given samples $x\sim q$ where $q$ represents the old policy and $p$ the new policy, the traditional unbiased estimator $k_{1}$ takes the form:
$k_{1}=-\log\left(\frac{p(x)}{q(x)}\right)$ (6)
While $k_{1}$ is unbiased, it exhibits high variance, which can potentially destabilize training. To address the high variance issue, an alternative estimator $k_{2}$ takes the form:
$k_{2}=\frac{1}{2}\left(\log\left(\frac{p(x)}{q(x)}\right)\right)^{2}$ (7)
This quadratic formulation has lower variance but introduces bias into the estimation. To achieve both unbiasedness and low variance, a third estimator $k_{3}$ takes the form:
$k_{3}=(r-1)-\log(r)$ (8)
where $r=p(x)/q(x)$ denotes the density ratio. In practice, policy networks output log probabilities, and thus $\log r$ is directly computed as the difference between log probabilities, i.e., $\log r=\log p(x)-\log q(x)$. $k_{3}$ theoretically achieves a balance between variance and bias while maintaining the non-negativity of KL divergence. In this work, we adopt a $k_{3}$ estimator following recent practice of GRPO (Shao et al., 2024).

### v6

@@ Abstract
Xueru Wen1,2, Jie Lou3, Xinyu Lu1,2, Ji Yuqiu3, Xinyan Guan1,2, Yaojie Lu1∗, Hongyu Lin1, Ben He1,2∗, Xianpei Han1,2, Debing Zhang3, Le Sun1,2 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences, Beijing, China 2University of Chinese Academy of Sciences, Beijing, China 3Xiaohongshu Inc {wenxueru2022,luxinyu2021,guanxinyan2022}@iscas.ac.cn {luyaojie,hongyu,sunle,xianpei}@iscas.ac.cn benhe@ucas.edu.cn loujie0822@gmail.com dengyang@xiaohongshu.com
footnotetext: *Corresponding authors.
@@ 1 Introduction
    Comprehensive evaluations on HotpotQA, SQuADv2, and Biography present significant improvements of RLFH over both base models and existing hallucination mitigation approaches, showing the method’s effectiveness [^1] The code and data associated with this work are available at https://github.com/AlignRM/RLFH..
@@ 3.2.2 Online Reinforcement Learning
Specifically, we first sample the prompt $x$ and corresponding response $y$. Subsequently, the policy model $\pi$ itself serves as the judge to provide fine-grained feedback through the assessment framework. This feedback is then converted into token-level dense reward $r=\left[r_{1},r_{2},...,r_{T}\right]$, where $T$ denotes the total length of the response $y$. Finally, we employ this reward $r$ to update the policy model $\pi$ by the Proximal Policy Optimization (Schulman et al., 2017) algorithm.
@@ 4.2 Main Results
Table 2: Results of baseline and RLFH-trained models using different granularity reward signals.
@@ 5 Conclusion
Acknowledgment
We sincerely thank the reviewers for their insightful comments and valuable suggestions. This work was supported by Beijing Natural Science Foundation (L243006), Beijing Municipal Science and Technology Project (Nos. Z231100010323002), the Natural Science Foundation of China (No. 62306303, 62272439, 62476265) and the Basic Research Program of ISCAS (Grant No. ISCAS-JCZD-202303).